# ElectricityEmissions.jl: A Framework for the Comparison of Carbon Intensity Signals

### v2

@@ Abstract.
An increasing number of individuals, companies and organizations are interested in computing and minimizing the carbon emissions associated with their real-time electricity consumption. To achieve this, they require a carbon signal, i.e. a metric that defines the real-time carbon intensity of their electricity supply. Unfortunately, in a grid with multiple generation sources and multiple consumers, there is no unambiguous way to trace electricity from source to sink. This makes it hard to define an appropriate signal, leading to a raging discussion about how to best quantify the carbon footprint of electricity. This paper seeks to inform the discussion about which carbon signal is better or more suitable for two important use cases, namely carbon-informed load shifting and carbon accounting. We do this by developing a new software package ElectricityEmissions.jl, that computes several established and newly proposed carbon emission metrics for standard electric grid test cases. We also demonstrate how the package can be used to investigate the effects of using these metrics to guide load shifting. Our results affirm previous research, which showed that the choice of carbon emission metric has significant impact on shifting results and associated carbon emission reductions. In addition, we demonstrate the impact of load shifting on both the consumers that perform the shifting and consumers that do not. Disconcertingly, we observe that shifting according to common metrics such as average carbon emissions can reduce the amount of emissions allocated to data center, but cause an increase in the total emissions of the system.
@@ 1. Introduction
As the impacts on climate change become increasingly visible, many electricity consumers, from private persons to large corporations, show increasing interest in assessing and reducing their carbon footprint. For many consumers, a significant portion of their carbon footprint is tied to their electricity consumption and associated emissions from electricity generation. The source of electricity generation (e.g. solar, wind, hydro, nuclear, natural gas or coal) as well as the temporal availability of renewable power (i.e. daily and seasonal cycles of solar, wind and hydropower) vary widely between locations, leading to differences in carbon intensity of electric power across time and space. This provides an opportunity for consumers to reduce carbon emissions by using power where and when low-carbon electricity is available. However, to actively reduce the carbon emissions of their electricity supply, consumers need real-time information about how carbon emissions of electricity vary. Furthermore, they need a framework for leveraging this information not only for real-time load shifting but also for long-term carbon accounting, such that they can get credit for their efforts to reduce emissions. This is particularly important for companies and industries that are subject to emerging carbon emission regulations and reporting requirements, including e.g. producers of clean hydrogen (Internal Revenue Service, 2023).
Quantifying the carbon intensity of electricity consumption at differing times/grid locations, however, is not straightforward. While it is possible obtain information on emissions from the generators and assess how the total carbon footprint varies over time, the physics of the grid do not lend themselves to a clear definition of generation mix (and hence emissions intensity) at individual locations. This leaves significant confusion and open questions regarding what the best metric for carbon emissions, which is the topic of ongoing debate (Sukprasert et al., 2024). This paper develops a software package and an analysis framework to help evaluate the effectiveness of different carbon emission metrics.
@@ 1.1. Related work
Past research has focused on computing and analyzing different carbon emission metrics for real-time electricity consumption (e.g. (Electricitymaps, [n. d.]; Callaway et al., 2018; WattTime, [n. d.]; Lindberg et al., 2021b; Chen et al., 2023b, c)) or quantified impacts of current standards for long-term carbon accounting on long-term emissions (see (Langer et al., 2023) and references therein). Recently, there has also been increased interest in comparing these metrics for the purposes of carbon accounting and informing load-shifting applications. (Sukprasert et al., 2024; Richardson, 2023; Corradi, 2023)
The two main shifting metrics that are currently in use and being debated are ‘Average Carbon Emissions’ (ACE), defined as total system emissions divided by total system load, and ‘Locational Marginal Carbon Emissions’ (LMCE), defined as the emissions factor of the marginal generator(s) (i.e. the generator(s) that would respond if a load changes its consumption by a small amount). In a recent paper, the authors of (Sukprasert et al., 2024) compare marginal and average carbon emissions signals for the purpose of data-center workload scheduling. Their results, based on data from 65 regions throughout the world, found the two signals to be largely negatively-correlated, leading to significant differences in workload scheduling outcomes as well as accounted emissions.
@@
Inter-metric variability in shifting-incentive and emissions accounting properties, such as that demonstrated in (Sukprasert et al., 2024), highlight the need for new tools and analysis frameworks to compare different emissions metrics. This is particularly important given that ACE and LMCE both have their drawbacks, which has led to interest in defining new metrics such as those based on ‘carbon flow’ methods (Chen et al., 2023a, b).
@@
To inform the debate about the effectiveness and appropriateness of different carbon emission metrics, we develop a software package for comparing the different metrics from a power system perspective. Our comparison framework is based on emulating the market clearing mechanism of an electric grid (represented by solving a DC optimal power flow (OPF)) and computing total carbon emissions for the overall grid as well as various carbon emission metrics for different locations based on the result. This information can then be used for both carbon accounting (i.e. to allocate carbon emissions to different loads) and as an input to simulate real-time load-shifting, where a carbon-sensitive load, such as a data center, changes their consumption in response to the carbon metrics. Finally, it is possible to assess the effectiveness of real-time load shifting by re-running the simulation of the electricity market clearing with the updated electricity consumption pattern.
@@
One important benefit of the framework is that it allows us (and other researchers) to evaluate both whether a given metric is effective in guiding real-time load shifting (i.e. if a consumer reduces their use of electricity in an hour with a high emission value and increases it in an hour with low emission value, does this lead to a reduction in overall system emissions?) and whether it possesses desirable properties for carbon accounting (e.g. does the metric guarantee that the total carbon emissions assigned to consumers are equal to the total carbon emissions from generation?). This can help inform the debate around current metrics. However, our second goal is to provide a platform for evaluation of newly proposed metrics. We demonstrate this by proposing a new carbon emission metric, namely the adjusted locational marginal emissions (ALMCE) metric, and assessing its merits relative to existing metrics using our framework.
@@ 2. Background
We consider an idealized setting, where the market clearing can be modelled mathematically as an optimal power flow (OPF) problem. The OPF problem minimizes the overall cost of operating the electric grid, subject to constraints that model the physical flows of electric power and limits on generation and transmission capacity. It can be expressed as a linear optimization problem. We provide the mathematical formulation for this optimization problem in Appendix A.
The OPF problem is solved approximately every 5-15 min (i.e. every time step $t$) in real-time operations of the electric grid. The main outcome of the OPF is the optimal generation dispatch, given by a set of generation setpoints $p_{G,t}^{*}$. The reason for repeatedly solving the OPF is to update the optimal dispatch $p_{G,t}^{*}$ in response to changes in the power system load $p_{D,t}$ and other system parameters, such as the time-varying availability of renewable generation represented by the time-varying maximum generation limits $p_{G,t}^{max}$.
@@
In the OPF problem, increasing or decreasing consumption $p_{D,t+1}$ will change the optimal generation set-points $p_{G,t+1}^{*}$ and the resulting carbon footprint of the grid. This allows consumers to impact carbon emissions by changing their consumption, for example by shifting their consumption to a different location (e.g. by shifting computing loads within networks of data centers) or to a different time step (e.g. by deferring consumption until later in the day).
@@ 3. Carbon Emission Metrics
To define the carbon emissions of electricity consumption, we need to define a time- and location-specific carbon emission metric $e_{D,j,t}$ at each location in the grid, such that the carbon emissions of load $j$ at time $t$ given by
@@
The challenge lies in choosing how to define $e_{D,j,t}$ (often also referred to as carbon intensity or emissions factor). Even though the carbon emission factor from each electric generator $i$ and time step $t$, denoted by $e_{G,i,t}$, are well defined, there is no clear consensus about how to allocate these carbon emissions to different consumers. This is in part because there is no unambiguous method to trace the actual flow of electricity from generator to load. Thus, multiple definitions of carbon emission metrics have been proposed. We summarize the definitions of some widely used metrics below, along with more recently proposed one and a brand new metric that we consider here for the first time. Note that this is not intended to be an exhaustive list of carbon metrics. Some other metrics, such as long-run marginal emissions (Hawkes, 2014), are not included below as they are less relevant to real-time carbon accounting and load shifting.
@@
The most intuitive metric is the average carbon emissions $e_{D,t}^{avg}$, defined as the total carbon emissions $E_{tot,t}$ and divided by the total power consumption,
(3) $e_{D,j,t}^{avg}=e_{D,t}^{avg}=\textstyle{\frac{E^{tot}_{G,t}}{\sum_{j\in\mathcal{D}}p_{D,j,t}}}.$
With this definition, the emissions factor $e_{D,j,t}$ is the same for all loads $j$. As a metric, average carbon emissions has several benefits. It is easily accessible (i.e. published by system operators or estimated from publicly available information), is available for large geographical regions from data providers such as ElectricityMaps (Electricitymaps, [n. d.]) and WattTime (WattTime, [n. d.]), and can be used for scope 2 electricity accounting under the Greenhouse Gas Protocol (GHG Protocol, 2015). Average carbon emissions capture the emissions from all generators, and ensures that the total emissions from generation equal the total emissions assigned to loads. However, the average carbon emissions metric also has several drawbacks. It does not account for grid constraints that limit the transfer of power from one region to another. Further, average carbon emissions fails to capture how load shifting, i.e. changes in the timing or location of electricity consumption $p_{D,t}$, impacts total grid emissions, and does not reflect the availability of low-carbon generation sources, and, in particular, the presence of curtailed renewable energy.
@@
We illustrate these drawbacks based on an example from the CAISO grid (in California) on June 10, 2021. Fig. 1 shows the total generation for each fuel type (top), average carbon emissions measured in gCO2eq/MWh (middle) and amount of renewable curtailment in the grid (bottom). We observe that the average emissions are lowest in the late afternoon, when the renewable energy generation is highest. However, there is significant renewable energy curtailment earlier in the day, which is why the average emissions are higher in this time period. If the average emissions metric is used to shift load, consumers may choose to wait until later in the day with using power. However, from a grid perspective, this is opposite of what is needed, as such a shift would likely increase renewable energy curtailment and lead to overall higher emissions.
@@
To answer questions such as “how do carbon emissions change if I choose to charge my electric vehicle now rather than later?”, we can instead consider the marginal carbon emissions. By applying sensitivity analysis methods to the OPF problem, we can describe how a small change in a given load $\delta p_{D,j,t}$ changes the optimal generation dispatch $\delta p_{G,t}^{*}$. This allows us derive the associated change in emissions $\delta E=e_{G,t}^{T}\delta p_{G,t}^{*}$ and obtain the so called locational marginal carbon emissions (LMCE)
(4) $e_{D,j,t}^{lmce}=\textstyle{\frac{\delta E}{\delta p_{D,j,t}}}.$
@@
Our implementation of LMCE is based on the method described in (Lindberg et al., 2021b), but extended to account for piece-wise linear cost functions. The full mathematical model is described in Appendix B. Note that LMCE inherently account for transmission constraints, and this varies by location in the grid when transmission capacity is limited.
@@
The locational marginal emissions $\lambda_{CO2}$ are arguably the metric that most accurately describes the impact of (small) load shifting actions (i.e. intentional changes in $p_{D,t}$) to reduce carbon emissions (Lindberg et al., 2021a), (Lindberg et al., 2022). However, because they represent sensitivities of the OPF solution, LMCE can vary widely between time-steps and and may not be an appropriate metric for larger load shifts.
@@
Further, it is unclear how LMCE can be used for carbon accounting, as it does not capture any information about emissions from non-marginal generators. Here, it is important to note the distinction between LMCE and locational marginal prices (LMPs), which are used to clear the electricity market. Since electricity markets seek to minimize total generation cost, the locational marginal price represent the marginal generation cost of the most expensive generator (i.e. the last one to get dispatched to serve electric load at a given location). Other generators that are running have a lower cost. For emissions, there is no such relationship, i.e. the marginal generator may have higher or lower emissions than the other units that are dispatched. As a result, if we allocate carbon emissions to loads based on the locational marginal emissions, the total allocated carbon emissions will generally be higher or lower than the total emissions from generation. In many cases we find that the difference between the total emissions from generation is significantly different from the total emissions allocated to loads.
@@
To illustrate the volatility of LMCE, Fig. 2 shows the average and locational marginal emissions for ComEd in Chicago on 8/13/23. We observe that the 5 minute, real-time locational marginal emissions have much higher volatility than the hourly average emissions. We also see that the LMCE is significantly higher than the average emissions, indicating that the marginal generator at this location is more polluting than the average generator.
@@
Another class of metrics that have been considered in the academic literature are metrics based on a “carbon flow” which “traces” carbon emissions from generators to load (Kang et al., 2015; Chen et al., 2023c, b). The key idea behind these metrics is that generators inject both power and carbon emissions into the grid, and the carbon “flows” along with the electricity. The methods in (Kang et al., 2015; Chen et al., 2023c, b) are based on the so-called proportional power sharing rule (Bialek, 1996; Kirschen and Strbac, 1999), an assumption which stipulates that the power flowsinto a bus (and the carbon emissions flowing with it) are shared proportionally across the power flows out of it. It is important to note that this an assumption, not a physical law. There exist a variety of other proposals for tracing power flow (see e.g. (Chen and Dhople, 2019) and references therein), which would likely lead to different carbon allocations across loads. To the knowledge of the authors, carbon flow metrics have not been adopted in practice, possibly because their definition is less intuitive and more mathematically involved, and because the computation requires detailed knowledge regarding the network topology and flow of power. Nevertheless, these methods are interesting as they give rise to a range of possible carbon emission metrics, including locational average carbon emissions (Chen et al., 2023b).
@@
(5) $e_{D,j,t}^{ame}=e_{D,j,t}^{lme}+\frac{E_{G,t}^{tot}-E_{D,t}^{lme}}{\sum_{j\in\mathcal{D}}p_{D,j,t}},$
where the first term is the locational marginal emissions $e_{D,j,t}^{lme}$ and second term represents the adjustment. This adjustment is defined as the difference between the total emissions from generation $E_{G,t}^{tot}$ and the total emissions assigned to loads using the LMCE metric $E_{D,t}^{lmce}=\sum_{j\in\mathcal{D}}e_{D,j,t}^{lmce}p_{D,j,t}$, shared across all loads proportional to their consumption.
While the full properties of ALMCE for carbon accounting and load-shifting remain to be determined through further analysis empirical testing, there are some properties worth noting. First, if there is a single marginal generator in the system, ALMCE is equivalent to ACE and in this case would lead to similar shifting results as ACE. Second, if we consider a single time step and sort nodes according to their nodal carbon emission values (e.g. from high to low), then ALMCE and LMCE would lead to the same ordering of nodes. Therefore, if we only consider spatial load shifting (where you can choose between consuming at different locations rather than at different times), ALMCE and LMCE might lead to similar shifting results.
@@
We have created a new Julia package, ElectricityEmissions.jl  [^1] Link to package, removed for doubleblind review , to serve as a central implementation of the carbon metrics mentioned in the previous section. This package is open-source, and contributions of other carbon metrics and accounting methods are welcome. Below is an example of computing and plotting carbon metrics for a small five-bus network.

### v3

@@ Abstract.
An increasing number of individuals, companies and organizations are interested in computing and minimizing the carbon emissions associated with their real-time electricity consumption. To achieve this, they require a carbon signal, i.e. a metric that defines the real-time carbon intensity of their electricity supply. Unfortunately, in a grid with multiple generation sources and multiple consumers, the physics of the system do not provide an unambiguous way to trace electricity from source to sink. As a result, there are a multitude of proposed carbon signals, each of which has a distinct set of properties and method of calculation. It remains unclear which signal best quantifies the carbon footprint of electricity. This paper seeks to inform the discussion about which carbon signal is better or more suitable for two important use cases, namely carbon-informed load shifting and carbon accounting. We do this by developing a new software package ElectricityEmissions.jl, that computes several established and newly proposed carbon emission metrics for standard electric grid test cases. We also demonstrate how the package can be used to investigate the effects of using these metrics to guide load shifting. Our results affirm previous research, which showed that the choice of carbon emission metric has significant impact on shifting results and associated carbon emission reductions. In addition, we demonstrate the impact of load shifting on both the consumers that perform the shifting and consumers that do not. Disconcertingly, we observe that shifting according to common metrics such as average carbon emissions can reduce the amount of emissions allocated to the consumer doing the shifting, while increasing the total emissions of the power system.
@@ 1. Introduction
As the impacts of climate change become increasingly visible, many electricity consumers, from private persons to large corporations, have shown increasing interest in assessing and reducing their carbon footprint. For many consumers, a significant portion of their carbon footprint is tied to their electricity consumption and associated emissions from electricity generation. The source of electricity generation (e.g. solar, wind, hydro, nuclear, natural gas or coal) as well as the temporal availability of renewable power (i.e. daily and seasonal cycles of solar, wind and hydropower) vary widely between locations, leading to differences in carbon intensity of electric power across time and space. This provides an opportunity for consumers to reduce carbon emissions by using power where and when low-carbon electricity is available. However, to actively reduce the carbon emissions of their electricity supply, consumers need real-time information about how carbon emissions of electricity vary. Furthermore, they need a framework for leveraging this information not only for real-time load shifting but also for long-term carbon accounting, such that they can get credit for their efforts to reduce emissions. This is particularly important for companies and industries that are subject to emerging carbon emission regulations and reporting requirements, including e.g. producers of clean hydrogen (Internal Revenue Service, 2023).
Quantifying the carbon intensity of electricity consumption in a grid with multiple consumers and multiple sources of generation, is however not straightforward. While it is possible to obtain information on emissions from the generators and assess how the total carbon footprint varies over time, the physics of the grid do not lend themselves to a clear definition of the generation mix (and hence emissions intensity) that serves individual locations. This leaves significant open questions regarding which carbon emissions is the best, see e.g. (Sukprasert et al., 2024; Ricks et al., 2024). This paper develops a software package and an analysis framework to help evaluate the effectiveness of different carbon emission metrics.
@@ 1.1. Related work
Existing research has focused on computing and analyzing different carbon emission metrics for real-time electricity consumption, e.g. in (Electricitymaps, [n. d.]; Callaway et al., 2018; WattTime, [n. d.]; Lindberg et al., 2021b; Chen et al., 2023b, c), spurring an interest in how consumers might use such emissions metrics to assess and minimize the carbon impact of their electricity consumption. Previous studies have examined the potential to reduce electricity emissions in the context of data-centers (Lin et al., 2012; Lindberg et al., 2021b; Zheng et al., 2020; Radovanovic et al., 2022), production of hydrogen fuel (Ricks et al., 2022; Zhang et al., 2020), and residential consumption (Mata et al., 2020; Harris et al., 2015). The increasing availability of real-world carbon emissions data, provided both directly by grid operators (PJM, 2024; National Energy System Operator, [n. d.]) and by third party organizations (WattTime, [n. d.]; Electricitymaps, [n. d.]), has enabled real-world implementations of load shifting strategies guided by emission metrics(Radovanovic et al., 2022; Maji et al., 2023; Inc., 2023; National Grid, [n. d.]). For example, battery-powered devices such as phones have settings that enable “clean energy charging” (Inc., 2023), while customers in UK can install light bulbs that turn green when the electricity is green (National Grid, [n. d.]).
Generally, existing works have tended to focus on a single carbon metric–e.g. to analyze its properties or use it as input to a theoretical or realized emissions-reduction scheme. Recently, however, there has also been increased interest in comparing different metrics for the purposes of carbon accounting and informing load-shifting applications (Lindberg et al., 2021a, 2022; Richardson, 2023; Corradi, 2023). Two main shifting metrics that are currently in use. ‘Average Carbon Emissions’ (ACE), defined as total system emissions divided by total system load, is the carbon metric used by the Greenhouse Gas Protocol for so-called location-based carbon accounting (GHG Protocol, 2015). ‘Locational Marginal Carbon Emissions’ (LMCE), defined as the change in emissions due to a small change in load, has been investigated in a number of research studies and is available from at least one grid operator. A variety of works have examined the effectiveness of ACE and LMCE for informing load-shifting applications (Lindberg et al., 2021a, 2022; Sukprasert et al., 2024). In (Lindberg et al., 2021a), a synthetic grid case study is used to demonstrate that the emissions impact of data-center load shifting based on either ACE or LMCE (as well as alternate data such as price and renewable curtailment) has a significant impact on shifting outcomes. In a recent paper, the authors of (Sukprasert et al., 2024) compare real-world marginal and average carbon emissions signals from 65 regions throughout the world, and found the two signals to be largely negatively-correlated, leading to significant differences in load shifting outcomes and accounted emissions.
The debate between ACE and LMCE, and resulting illumination of their respective shortcomings, has also led to interest in defining new classes of carbon metrics. One notable example of this are so-called ‘carbon-flow’-based metrics (Kang et al., 2015; Chen et al., 2023a, b), which employ various non-physical assumptions to trace power flow (and associated carbon emissions) from generators to loads. Though not yet extensively analyzed, the method by which these carbon flow metrics are calculated is distinct from both ACE and LMCE. As a result, such metrics likely represent yet more competing definitions of carbon intensity, highlighting the increasing need to study the pros and cons of each type of metric.
@@
Inter-metric variability in shifting-incentive and emissions accounting properties, such as that demonstrated in (Lindberg et al., 2021a) and (Sukprasert et al., 2024), highlights the need for new tools and analysis frameworks to compare different emissions metrics. This is particularly true given the high likelihood that new metrics will continue to be proposed and employed to inform consumers’ beliefs regarding their carbon footprint, as well as their load-shifting behavior.
@@
To inform the debate about the effectiveness and appropriateness of different carbon emission metrics, we develop a software package for comparing the different metrics from a power system perspective. Our comparison framework is based on modeling the US market clearing mechanism of an electric grid (represented by solving a DC optimal power flow (OPF) to determine the optimal cost-minimizing generator set-points) and computing total carbon emissions for the overall grid as well as various carbon emission metrics for different locations based on the result. This information can then be used for both carbon accounting (i.e. to allocate carbon emissions to different loads) and as an input to simulate load-shifting, where a carbon-sensitive load, such as a data center, changes their consumption in response to the carbon metrics. Finally, it is possible to assess the effectiveness of load shifting by re-solving the electricity market clearing with the updated electricity consumption pattern.
@@
One important benefit of the framework is that it allows us (and other researchers) to evaluate both whether a given metric is effective in guiding load shifting (i.e. if a consumer reduces their use of electricity in an hour with a high emission value and increases it in an hour with low emission value, does this lead to a reduction in overall system emissions?) and whether it possesses desirable properties for carbon accounting (e.g. does the metric guarantee that the total carbon emissions assigned to consumers are equal to the total carbon emissions from generation?). This can help inform the debate around current metrics. However, our second goal is to provide a platform for evaluation of newly proposed metrics. We demonstrate this by proposing a new carbon emission metric, namely the adjusted locational marginal emissions (ALMCE) metric, and assessing its merits relative to existing metrics using our framework.
@@ 2. Background
Notation
In the following discussions, we will denote quantities related to electricity consumption by subscript $D$ and the set of loads by $\mathcal{D}$, while quantities associated with electricity generation are denoted by subscript $G$ and the set of generators is denoted by $\mathcal{G}$. We will generally use lower case $e$ to refer to carbon emissions intensities (measured in [TonsCO2/MW]) and capital $E$ to refer to carbon emissions (measured in [TonsCO2]). Lower case $p$ will be used to denote quantities of power generation or consumption.
@@
We consider a region where the electricity market clearing can be modeled mathematically as an optimal power flow (OPF) problem. This is representative of an electricity market based on nodal pricing, which is the prevalent market design in the United States. In this market design, generators provide offers to generate electricity, which are associated with a given cost [\$/MWh] and capacity [MW]. The electricity consumption of electric loads is typically forecasted by the utilities that serve those loads, and the consumption is treated as a fixed value. This modeling is appropriate for most electric loads, which buy electricity on fixed rates through their local utility. However, the OPF model could easily be extended to also consider flexible, price-sensitive loads that source electricity directly from the whole sale energy market by modeling them as negative generation.
Given that the load is assumed to be fixed, the goal of the market clearing, and the objective of the corresponding OPF problem, is to minimize the total cost of generating electricity. This cost minimization is subject to constraints that model the physical flows of electric power and limits on generation and transmission capacity. It can be expressed as a linear optimization problem. We provide the mathematical formulation and additional explanations for this optimization problem in Appendix A.1.
Electricity markets clear in two periods, day-ahead (DA) and real-time (RT). The DA market, solves a multi-time period optimization problem for the generation at each hour of the following day. The RT market is solved every 5 minutes during the day, where discrepancies from the load forecast or renewable energy forecast used in the DA market are resolved by solving an OPF and adjusting the generator output. For either market, the primary outcome of the OPF is the optimal generation dispatch, given by a set of generation set-points $p_{G,i,t}^{*}$ for all generators $i\in\mathcal{G}$ at time step $t$, that is needed to satisfy system load $p_{D,j,t}$, representing the demand by loads $j\in\mathcal{D}$ at time $t$.
@@
In the OPF problem, increasing or decreasing consumption $p_{D,t+1}$ will change the optimal generation set-points $p_{G,t+1}^{*}$ and the resulting carbon footprint of the grid. This allows consumers to impact carbon emissions by changing their consumption, for example by shifting their consumption to a different location (e.g. by shifting computing loads within networks of data centers) or to a different time step (e.g. by deferring consumption until later in the day). We note that while it is possible that a consumer might shift their load to better align with behind-the-meter renewable assets (hence lowering their emissions), the OPF-based framework presented in this work considers only consumer impact on utility-scale generation.
@@ 3. Carbon Emission Metrics
To define the carbon emissions of electricity consumption, we need to define a time- and location-specific carbon emission metric $e_{D,j,t}$ [TonsCO2/MW] that defines the carbon intensity per unit of electricity consumed at each timestep and location in the grid. Given the carbon emission metric $e_{D,j,t}$, the total carbon emissions of demand $j$ at time $t$ are given by
@@
where $p_{D,j,t}$ [MW] denotes the total electricity consumption and $E_{D,j,t}$ [TonsCO2] denotes the total emissions. The challenge lies in choosing how to define $e_{D,j,t}$ (often also referred to as carbon intensity or emissions factor). Even though the carbon emission factor from each electric generator $i$ and time step $t$, denoted by $e_{G,i,t}$, are well defined, there is no clear consensus about how to allocate these carbon emissions to different consumers. This is in part because there is no unambiguous method to trace the actual flow of electricity from generator to load. Thus, multiple definitions of carbon emission metrics have been proposed. We summarize the definitions of some widely used metrics below, along with a more recently proposed metric and a brand new metric that we consider here for the first time. Note that this is not intended to be an exhaustive list of carbon metrics. Some other metrics, such as long-run marginal emissions (Hawkes, 2014), are not included below as they are less relevant to real-time carbon accounting and load shifting.
@@
The most intuitive and commonly used metric is average carbon emissions $e_{D,t}^{avg}$, which is defined as the total carbon emissions incurred by the generators $E_{G,t}^{tot}$ and divided by the total power consumption [^1] Note that care must be taken to appropropriately count the total carbon emissions from generation and the total load in cases where a region is importing or exporting electricity.,
(3) $e_{D,t}^{avg}=\textstyle{\frac{E^{tot}_{G,t}}{\sum_{j\in\mathcal{D}}p_{D,j,t}}}.$
With this definition, the emissions factor $e_{D,j,t}=e_{D,t}^{avg}$ is the same for all loads $j$ in the grid at time $t$. As a metric, average carbon emissions has several benefits. It is easily accessible (i.e. published by system operators or can be estimated from publicly available information without requiring a model of the network), is available for large geographical regions from data providers such as ElectricityMaps (Electricitymaps, [n. d.]) and WattTime (WattTime, [n. d.]), and can be used for scope 2 electricity accounting under the Greenhouse Gas Protocol (GHG Protocol, 2015). Average carbon emissions capture the emissions from all generators, and ensures that the total emissions from generation equal the total emissions assigned to loads. However, the average carbon emissions metric also has several drawbacks. It does not account for grid constraints that limit the transfer of power within a given region. Further, average carbon emissions does not accurately capture how load shifting, i.e. changes in the timing or location of electricity consumption of a specific load $p_{D,j,t}$, impacts total grid emissions. Another emerging drawback is that average carbon emissions does not reflect the availability of low-carbon generation sources that are currently not being used, and, in particular, the availability of curtailed renewable energy.
@@
We illustrate these drawbacks based on an example from the CAISO grid (in California) on June 10, 2021. Fig. 1 shows the total generation for each fuel type (top), average carbon emissions measured in TonsCO2/MWh (middle) and amount of renewable curtailment in the grid (bottom). We observe that the average emissions are lowest in the late afternoon, when the renewable energy generation is highest. However, there is significant renewable energy curtailment earlier in the day, which is why the average emissions are higher in this time period. If the average emissions metric is used to shift load, consumers may choose to shift their power consumption until later in the day. However, from a grid perspective, this is opposite of what is needed, as such a shift would likely increase renewable energy curtailment and lead to overall higher emissions.
@@
To answer questions such as “how do carbon emissions change if I choose to charge my electric vehicle now rather than later?”, a more accurate metric is the locational marginal carbon emissions. By applying sensitivity analysis methods to the OPF problem, we can describe how a small change in a given load, denoted by $\delta p_{D,j,t}$, changes the optimal generation dispatch. We denote changes in generation by $\delta p_{G,j,t}^{*}$. Once we know $\delta p_{G,j,t}^{*}$, we can derive the associated change in grid emissions at time $t$ as $\delta E_{G,t}^{tot}=\sum_{i\in\mathcal{G}}e_{G,i,t}\delta p_{G,i,t}^{*}$. After normalizing for the change in load, we obtain the so called locational marginal carbon emissions (LMCE)
(4) $e_{D,j,t}^{lmce}=\textstyle{\frac{\delta E_{G,t}^{tot}}{\delta p_{D,j,t}}}.$
@@
Our implementation of LMCE is based on the method described in (Lindberg et al., 2021b), but extended to account for piece-wise linear cost functions. The full mathematical model is described in Appendix A.2. Note that LMCE inherently accounts for transmission constraints, and thus varies by location in the grid when transmission capacity is limited.
@@
The locational marginal emissions $\lambda_{CO2}$ are arguably the metric that most accurately describes the impact of (small) load shifting actions (i.e. intentional changes in $p_{D,j,t}$) to reduce carbon emissions (Lindberg et al., 2021a), (Lindberg et al., 2022). However, because they represent sensitivities of the OPF solution, LMCE can vary widely between time-steps and and may not be an appropriate metric for larger load shifts.
@@
Further, it is unclear how LMCE can be used for carbon accounting, as it does not capture any information about emissions from non-marginal generators. Here, it is important to note the distinction between LMCE and locational marginal prices (LMPs), which are the result of clearing the electricity market. Since electricity markets seek to minimize total generation cost, the locational marginal price represents the marginal generation cost of the most expensive generator that is dispatched in the market clearing to serve electric load at a given location). The marginal generator can vary for different locations based on the network constraints, but all other dispatched generators are lower cost. For emissions, there is no such relationship for the marginal generator being the highest emitting dispatched generator. The marginal generator may have higher or lower emissions than the other units that are dispatched. As a result, if we allocate carbon emissions to loads based on the locational marginal emissions as the emission intensity, the total allocated carbon emissions will generally be higher or lower than the total emissions from generation. In many cases we find that the difference between the total emissions from generation is significantly different from the total emissions allocated to loads.
@@
To illustrate the volatility of LMCE, Fig. 2 shows the average and locational marginal emissions for ComEd in Chicago on 8/13/23. This is real-world data reported by the utility. We observe that the 5 minute, real-time locational marginal emissions have much higher volatility than the hourly average emissions. We also see that the LMCE is often significantly higher than the average emissions, indicating that the marginal generator at this location is more polluting than the average generator.
@@
Another class of metrics that have been considered in the academic literature are metrics based on a “carbon flow” which “traces” carbon emissions from generators to load (Kang et al., 2015; Chen et al., 2023c, b). The key idea behind these metrics is that generators inject both power and carbon emissions into the grid, and the carbon “flows” along with the electricity. The methods in (Kang et al., 2015; Chen et al., 2023c, b) are based on the so-called proportional power sharing rule (Bialek, 1996; Kirschen and Strbac, 1999), an assumption which stipulates that the power flows into a bus (and the associated carbon emissions) are shared proportionally across the power flows out of it. It is important to note that this is an assumption, not a physical law. There exist a variety of other proposals for tracing power flow (see e.g. (Chen and Dhople, 2019) and references therein), which would likely lead to different carbon allocations across loads. To the knowledge of the authors, carbon flow metrics have not been adopted in practice, possibly because their definition is less intuitive and more mathematically involved. Nevertheless, these methods are interesting as they give rise to a range of possible carbon emission metrics, including locational average carbon emissions (Chen et al., 2023b).
@@
(5) $e_{D,j,t}^{almce}=e_{D,j,t}^{lmce}+\frac{E_{G,t}^{tot}-E_{D,t}^{lmce}}{\sum_{j\in\mathcal{D}}p_{D,j,t}},$
where the first term is the locational marginal emissions $e_{D,j,t}^{lmce}$ and second term represents the adjustment. This adjustment is defined as the difference between the total emissions from generation $E_{G,t}^{tot}$ and the total emissions assigned to loads using the LMCE metric $E_{D,t}^{lmce}=\sum_{j\in\mathcal{D}}e_{D,j,t}^{lmce}p_{D,j,t}$, shared across all loads proportional to their consumption.
While the full properties of ALMCE for carbon accounting and load-shifting remain to be determined through further analysis empirical testing, there are some properties worth noting. First, ALMCE will fluctuate just as LMCE fluctuates, providing a rapid signal of the emissions of the marginal generators. Second,if there is a single marginal generator in the system, ALMCE is equivalent to ACE and in this case would lead to similar shifting results. Third, if we consider a single time step and sort nodes according to their nodal carbon emission values (e.g. from high to low), then ALMCE and LMCE would lead to the same ordering of nodes. Therefore, if we only consider spatial load shifting (where you can choose between consuming at different locations rather than at different times), ALMCE and LMCE might lead to similar shifting results. Finally, we note that the real-world implementation of ALMCE would not require significant effort in systems that already calculate LMCE, such as PJM (PJM, 2024), since the LMCE calculation is the most demanding part of computing ALMCE.
@@
We have created a new Julia package, ElectricityEmissions.jl  [^2] https://github.com/WISPO-POP/ElectricityEmissions.jl , to serve as a central implementation of the carbon metrics mentioned in the previous section. The package allows the calculation of these metrics for power system test-cases in the popular MATPOWER format, opening the door for carbon-intensity analysis on thousands of simulated grids of varying sizes and properties. ElectricityEmissions.jl is open-source, and contributions of other carbon metrics and accounting methods are welcome. Below is an example of computing and plotting carbon metrics for a small five-bus network.